OpenStudy (anonymous):

whats the final answer to: y+7=−7(x+1) Should be an equation for a tangent line*

OpenStudy (anonymous):

You know the equation for a tangent line, meaning you know how to find derivatives, but you're struggling with basic algebra?

OpenStudy (anonymous):

I got the results but for some reason its not matching up exactly with the possible answers.

OpenStudy (anonymous):

You're getting \(y=-7x-14\)?

OpenStudy (anonymous):

which is why i want someone to give me the answer so i can be sure i didn't make a mistake earlier in the problem.

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

What is f(x)?

OpenStudy (anonymous):

f(x) = -4x + 3/x at x = -1

OpenStudy (anonymous):

Yeah, I can tell already it's at the point (-1, -7) based on the form of your tangent line.

OpenStudy (anonymous):

Okay, so I'm getting: \[ f'(x) = -4 + -\frac{3}{x^2}\implies f'(-1) = -4-3 = -7 \]So that part is correct.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

You are fine so far.

OpenStudy (anonymous):

ok so what should i do next?

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

Write an equation for the tangent line to f(x) = -4x + 3/x at x=-1

OpenStudy (anonymous):

Oh wait, we need to find out \(\large f(-1)=-4(-1)+\frac{3}{-1}\)

OpenStudy (anonymous):

possible answers : A) -7x + y = -6 B) -6x + 2y = -5 C) 7x + y = -6 D) 6x + 2y = -5

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

4-3 = 1.... there was your problem.

OpenStudy (anonymous):

\[y-1=-7(x+1)\]

OpenStudy (anonymous):

Thank you for the help!